# Distribution Shifts at Scale: Out-of-distribution Detection in Earth Observation

### v1

@@ Abstract
Training robust deep learning models is critical in Earth Observation, where globally deployed models often face distribution shifts that degrade performance, especially in low-data regions. Out-of-distribution (OOD) detection addresses this challenge by identifying inputs that differ from in-distribution (ID) data. However, existing methods either assume access to OOD data or compromise primary task performance, making them unsuitable for real-world deployment. We propose TARDIS, a post-hoc OOD detection method for scalable geospatial deployments. The core novelty lies in generating surrogate labels by integrating information from ID data and unknown distributions, enabling OOD detection at scale. Our method takes a pre-trained model, ID data, and WILD samples, disentangling the latter into surrogate ID and surrogate OOD labels based on internal activations, and fits a binary classifier as an OOD detector. We validate TARDIS on EuroSAT and xBD datasets, across 17 experimental setups covering covariate and semantic shifts, showing that it performs close to the theoretical upper bound in assigning surrogate ID and OOD samples in 13 cases. To demonstrate scalability, we deploy TARDIS on the Fields of the World dataset, offering actionable insights into pre-trained model behavior for large-scale deployments. The code is publicly available at https://github.com/microsoft/geospatial-ood-detection
Figure 1: Illustration of the proposed method for OOD detection. Given a pre-trained model, a dataset of known ID samples, and WILD samples (containing both ID and OOD data), the method assigns surrogate ID or OOD labels to each WILD sample and fits a binary classifier $g$ (top row). During deployment, the classifier $g$ uses internal activations to predict whether an unseen sample is ID or OOD (bottom row).
@@ 1 Introduction
Deep learning (DL) models have demonstrated remarkable capabilities across various domains but often exhibit overconfidence in their predictions, even when confronted with data that diverges from their training distribution [25, 10, 19]. This overconfidence arises from the assumption that inference data will follow the same independent and identically distributed (i.i.d.) properties as the training data. However, in real-world applications, this closed-world assumption [8, 14] is frequently violated by test-time distribution shifts (e.g., lighting or angle variations, different devices or sensors) that can significantly degrade model performance and harm generalization. To address this, it is critical for predictive models to detect when new observations fall outside the training distribution, a task known as out-of-distribution (OOD) detection. Differentiating OOD from in-distribution (ID) samples is particularly challenging, partly due to the poor calibration of neural networks [6, 18]. This degradation poses a critical challenge, especially in applications where incorrect predictions can have severe consequences. In Earth Observation (EO), specifically with satellite imagery, distribution shifts are even more common due to variability in sensor types, acquisition angles, temporal changes, spatial variations, and landscape alterations caused by events such as natural disasters. This variability necessitates unique, tailored approaches for satellite imagery [21]. The nature of distribution shifts between training and inference data can be classified as near-distribution or far-distribution shifts, each posing unique challenges. For example, a model trained on data from a certain season and region may perform poorly when applied to other seasons in the same region or the same season in different regions [13], representing a far-distribution shift. Even models trained on data covering multiple seasons and regions may perform suboptimally under specific acquisition or atmospheric conditions, such as cloud cover, sensor faults, or artifacts, indicative of a near-distribution shift.
@@
Despite the critical importance of robust models in Earth Observation (EO), only a few studies address OOD detection in this domain [5, 3, 1]. These studies typically assume access to test-time distributions, rely on closed-set assumptions, and often sacrifice model performance on in-distribution tasks. This highlights the need for tailored OOD detection methods suited for large-scale EO deployment.
A further challenge arises when applying geospatial models globally. Many geospatial products, while measuring the same ground realities with high coverage, exhibit low levels of agreement [2]. Additionally, global geospatial models often suffer significant performance drops in low-data regions, such as sub-Saharan Africa [12]. These issues emphasize the need for global models equipped to detect OOD samples during inference.
OOD detection capabilities in real-world applications enhance model reliability by identifying when input data deviates from the training distribution. They also support resource allocation and targeted data collection by pinpointing areas where the model may underperform, effectively guiding both computational and human resources. Additionally, OOD detection enables real-time adjustments, allowing models to remain reliable in dynamic conditions. This approach supports a human-in-the-loop strategy for critical decisions, fostering trust, transparency, and accountability. Overall, OOD detection is essential for managing large-scale EO deployments effectively.
In this paper, we begin with an essential observation: OOD samples can trigger internal activation patterns that diverge from those of ID samples. Building on this insight and recognizing the challenges of deploying geospatial models at scale, we introduce the following key contributions:
  • We propose TARDIS (Test-time Addressing of Distribution Shifts at Scale), a post-hoc OOD detection method that preserves the model’s performance on the main task and operates without requiring access to unknown test-time near-distribution shifts.
  • We evaluate TARDIS on EuroSAT, a patch-classification dataset, and xBD, a semantic segmentation dataset, examining both covariate and semantic shifts across 16 experimental setups. For covariate shifts, we introduce geographical variance (spatial changes), temporal variance (seasonal changes), and environmental variance (landcover or disaster type changes). Semantic shifts are introduced by withholding a class during training and exclusively testing on that class, thus presenting new semantic categories.
  • We demonstrate that TARDIS can be deployed at scale, providing actionable insights into the trustworthiness and robustness of geospatial models in real-world scenarios.
@@ 2 Related Works
Our method lies at the intersection of two notable categories of OOD detection methods: scoring functions and activation manipulation. Scoring functions assign a numerical score to each input, reflecting its alignment with ID samples based on the model’s output. Maximum Softmax Probability (MSP) [10] detects OOD samples by assessing the softmax confidence score, assuming that low confidence indicates OOD samples. The Out-of-DIstribution detector for Neural networks (ODIN) [16] enhances MSP by applying input perturbations and temperature scaling to improve ID-OOD separation. The Mahalanobis score [15] calculates the distance between input features and class means in feature space, flagging inputs far from these means as OOD. The energy score [17] evaluates the model’s energy function to assess the likelihood of an input belonging to the ID distribution. Another category of OOD detection methods manipulates the internal activations of a pre-trained model to improve detection performance. ReAct [24] identifies differences in activation patterns of the penultimate layer between ID and OOD samples, enhancing separation by clipping activations at an upper limit. DICE [23] applies weight sparsification on a specific layer to further distinguish ID from OOD data and, when combined with ReAct, can enhance detection performance. Activation Shaping [4] prunes a portion of an input sample’s activations and slightly adjusts the remaining activations; when paired with the energy score, this approach has been shown to outperform contemporary OOD methods. Neuron Activation Patterns (NAP) [20] extracts, downsamples, and binarizes activation patterns from convolutional layers, computing the smallest Hamming distance between binarized test and training patterns. This distance provides a measure of model uncertainty, aiding in OOD detection.
Both scoring functions and activation-based strategies face notable challenges in dynamic, real-world environments. Scoring functions, while maintaining ID accuracy, struggle with OOD detection due to their reliance on static data distributions, an assumption rarely valid in practice. Activation-based methods, though effective, often compromise ID task performance and rely on hyperparameters tuned with hypothetical or substitute OOD samples. This reliance limits their effectiveness under evolving test-time conditions and shifting distributions.
@@
The task of OOD detection is inherently challenging, and the EO domain is no exception due to the diverse and heterogeneous nature of satellite imagery. Despite the critical importance of OOD detection in EO, this area remains relatively underexplored, with only a limited number of studies addressing the issue. One such study develops a Dirichlet Prior Network to quantify distributional uncertainty in DL models for satellite image analysis [5]. This method assumes Dirichlet distributions for ID samples and employs various setups where classes, color channels, and environmental features are alternately treated as ID or OOD. Another study frames anomaly detection as a cumulative open-set detection and location separation task [3]. This approach also uses a Dirichlet prior, similar to the first study, to expand the representation space between ID (normal) and OOD (anomalous) samples by predicting the anticipated categorical distribution. The method is suitable for scenarios where pre-event images are either unavailable, subject to radiation differences, or not sufficiently recent to aid in detection.
In this paper, we do not aim to optimize OOD detection performance but instead propose a method designed for real-world applications where the distribution is unknown, maintaining ID task performance is critical, and computationally expensive methods are impractical. Our focus is on handling near-distribution shifts, where data variations stem from the same satellite sources but exhibit changes over time, geography, or environmental factors. This contrasts with far-distribution shifts, where variations arise from entirely different data sources, such as different satellite models or natural images.
@@ 3 Problem Formulation
4 Method
Fitting a classifier to distinguish between in-distribution (ID) and out-of-distribution (OOD) samples typically requires known ID and OOD labels, which is impractical in real-world deployments. This creates a need for surrogate ID and OOD labels for samples with unknown distributions, which we refer to as WILD samples. Our method, illustrated in Figure 1, operates with a given model, a set of known ID samples, and a set of WILD samples. It generates surrogate ID and OOD labels for the WILD samples, which are then used to train a binary classifier that can classify the distribution of incoming samples.
First, we sample $M$ in-distribution samples from the ID set and $N$ WILD samples from the WILD set. Next, we pass both the ID and WILD samples through the frozen pre-trained model and extract activations from a specified layer. A pooling operation is applied to perform spatial downsampling, transforming the activations into one-dimensional vectors and combining the downsampled features. We then use k-means clustering to partition the combined features into $K$ clusters. For each cluster, we calculate the proportion of ID samples and assign surrogate labels based on a threshold $T$: if the proportion of ID samples in the cluster is greater than or equal to $T$, the surrogate label is 0; otherwise, it is 1.
Here, $y_{\text{sur}}(\mathbf{z})$ denotes the surrogate label: 00 for surrogate ID and $1$ for surrogate OOD. With this labeled dataset, we fit a binary classifier $g(\mathbf{z};\theta)$, which maps a sample to a predicted label. During deployment, for a sample of unknown distribution, we perform activation extraction and classification as follows: $\hat{y}=g\left(\text{Downsample}\left(f(\mathbf{x})\right);\theta\right)$. The output can be interpreted as probability scores, allowing it to quantify the magnitude of domain shift or thresholded to produce binary values (00 for in-distribution and $1$ for out-of-distribution).
Surrogate Label Assignment. The core novelty of TARDIS lies in its surrogate labeling generation step, which enables us to assign labels to WILD samples based on their proximity to ID samples in feature space. The assumption is that ID and OOD features are sufficiently distinct to allow effective separation through clustering. Specifically, we use k-means clustering to label WILD samples as surrogate OOD if WILD samples outnumber ID samples in a cluster, and as surrogate ID otherwise. Without access to true OOD data, ID samples serve as an indirect benchmark for distinguishing between ID and WILD samples. To optimize this process, we tune the number of clusters $K$ and threshold $T$ by balancing three objectives: entropy (to encourage homogeneity within clusters and distinct groups of surrogate labels), correct ID proportion (to maximize accurate identification of ID samples), and incorrect ID proportion (to minimize the misclassification of ID samples as OOD). The combined objective function minimizes average entropy and the misclassification rate of ID samples while maximizing correct ID classification, resulting in well-separated and accurately labeled groups.
Figure 2 illustrates our method. First, ID samples are drawn from the model’s training set, while WILD samples, representing unknown distributions, are sourced from satellite images taken at any location and time (Step 1). Both ID and WILD samples are passed through a pre-trained model, and internal activations are extracted from a specified layer (Step 2). Using these activations, we assign surrogate labels to WILD samples, categorizing them as surrogate ID or surrogate OOD (Step 3). This surrogate labeling, central to our approach, decomposes WILD samples based on their distributional alignment with ID samples. Finally, these labeled feature representations are used to train a binary classifier that, during deployment, outputs probability scores or thresholded 0/1 predictions to classify samples as either ID or OOD (Step 4).
@@ 5 Experimental Setup
In this section, we describe the experimental setups designed to evaluate TARDIS under controlled distribution shifts, which mimic real-world geospatial deployment challenges. Specifically, we use the EuroSAT [9] dataset for patch-level classification and the xBD [7] dataset for semantic segmentation. EuroSAT contains 27,000 labeled images across ten land use and land cover classes: Annual Crop, Forest, Herbaceous Vegetation, Highway, Industrial, Pasture, Permanent Crop, Residential, River, and Sea/Lake. The xBD dataset includes over 1 million labeled samples for building damage assessment from satellite imagery, covering multiple disaster types such as hurricanes, floods, and wildfires. For our experiments, we formulate xBD as a building footprint segmentation task with two classes (building footprint and background). EuroSAT and xBD provide a suitable test-bed for our OOD method due to their differences in volume, task, sensor type, geospatial distribution, and temporal and spatial characteristics, allowing a thorough evaluation under conditions commonly encountered in geospatial deployments.
@@
To introduce controlled distribution shifts, we reformulate the train, validation, and test splits of EuroSAT and xBD so that the model, when trained on the modified training set and evaluated on the adjusted test set, encounters shifts similar to those seen in geospatial model deployments. We categorize these shifts as covariate shift and semantic shift. Covariate shift occurs when the input data distribution changes between training and testing. This can introduce selection bias, as covariates in the training set may still appear in the test set to some extent, potentially skewing robustness evaluation. To address this, we also introduce semantic shift experiments where a class is held out during testing, creating an unseen semantic class. For our experiments, covariate shifts include spatial variance (varying spatial locations), temporal variance (varying time of data collection), and environmental variance (varying landcover or disaster types). For semantic shifts in the EuroSAT dataset, we train the model on nine out of ten classes and test on the holdout class, repeating this process so that each class serves as the holdout class once. Details of the experimental setups are provided in Table 1.
The setups described above assume clear definitions of ID (training set) and OOD (test set), which presupposes access to test-time distribution shifts, an assumption that is not feasible in real-world deployments. To evaluate our method’s performance under unknown test-time distribution shifts, we fit and evaluate the binary classifier $g$ in two distinct ways. In the first approach, we bypass surrogate label assignment (Step 3 in Figure 2), using the well-defined ID and OOD samples to fit and evaluate the classifier, referred to as $g_{\text{oracle}}$. In the second approach, we treat the OOD labels in the test set as unknown, simulating real-world conditions where these labels are unavailable. This method uses surrogate label assignment to reassign unknown labels into surrogate ID and surrogate OOD. Parameter estimation and evaluation on these reassigned labels yield a classifier denoted as $g^{*}$, representing our proposed method. The underlying assumption is that $g_{\text{oracle}}$ provides an upper bound on performance for $g^{*}$, as $g_{\text{oracle}}$ is trained with clearly separated ID and OOD samples, whereas $g^{*}$ is trained on surrogate ID and surrogate OOD samples that are generated through our surrogate label assignment process.
Figure 3: Geographical distribution of ID and WILD sets, containing 500 and 1200 samples, respectively. The ID set is sampled from the FTW dataset training set, while the WILD set is randomly sampled from the Microsoft Planetary Computer. Each Sentinel-2 patch, provided in two different time frames (planting and harvesting). The model $f$ takes both Sentinel-2 images from different seasons as input, and the predictions are shown on the right. The bar indicates the probability output of $g$ for each input pair, with a higher probability (closer to 1) suggesting a higher likelihood of the pair being OOD. On the map, probability values are thresholded at 0.5.

### v2

@@ Abstract
Training robust deep learning models is crucial in Earth Observation, where globally deployed models often face distribution shifts that degrade performance, especially in low-data regions. Out-of-distribution (OOD) detection addresses this by identifying inputs that deviate from in-distribution (ID) data. However, existing methods either assume access to OOD data or compromise primary task performance, limiting real-world use. We introduce TARDIS, a post-hoc OOD detection method designed for scalable geospatial deployment. Our core innovation lies in generating surrogate distribution labels by leveraging ID data within the feature space. TARDIS takes a pre-trained model, ID data, and data from an unknown distribution (WILD), separates WILD into surrogate ID and OOD labels based on internal activations, and trains a binary classifier to detect distribution shifts. We validate on EuroSAT and xBD across 17 setups covering covariate and semantic shifts, showing near-upper-bound surrogate labeling performance in 13 cases and matching the performance of top post-hoc activation- and scoring-based methods. Finally, deploying TARDIS on Fields of the World reveals actionable insights into pre-trained model behavior at scale. The code is available at https://github.com/microsoft/geospatial-ood-detection
Figure 1: Overview of the proposed OOD detection method. Given a pre-trained model, ID samples, and WILD samples (from unknown distributions), TARDIS assigns surrogate ID/OOD labels to WILD samples using the ID set and fits a binary classifier $g$ (top row). During deployment, $g$ uses internal activations of unseen samples to predict whether they are ID or OOD (bottom row).
@@ 1 Introduction
Deep learning models have demonstrated remarkable capabilities across various domains but often exhibit overconfidence in their predictions, even when confronted with data that diverges from their training distribution [31, 13, 24]. This overconfidence arises from the assumption that inference data will follow the same independent and identically distributed (i.i.d.) properties as the training data. However, in real-world applications, this closed-world assumption [11, 18] is frequently violated by test-time distribution shifts (e.g., lighting or angle variations, different devices or sensors) that can significantly degrade model performance and harm generalization. To address this, it is critical for predictive models to detect when new observations fall outside the training distribution, a task known as out-of-distribution (OOD) detection. Differentiating OOD from in-distribution (ID) samples is particularly challenging, partly due to the poor calibration of neural networks [9, 23]. This degradation poses a critical challenge, especially in applications where incorrect predictions can have severe consequences.
In Earth Observation (EO), specifically with satellite imagery, distribution shifts between training and inference commonly occur in the form of covariate and semantic shifts. Covariate shift refers to changes in the input distribution while keeping the task or label space fixed, whereas semantic shift involves changes in the label distribution or the meaning of inputs. A specific case of covariate shift is known as subpopulation shift [17], where the overall data distribution changes due to different proportions of underlying subgroups, even though the subgroups themselves remain the same. In EO, this arises when certain seasons, regions, or acquisition conditions are underrepresented during training but appear more prominently at test time. The magnitude of distribution shifts can further be characterized along a spectrum from near-distribution to far-distribution shifts. Near-distribution shifts may occur within the same satellite source due to temporal, geographic, or environmental variation—such as seasonal changes, sensor degradation, or atmospheric effects like cloud cover [16]. Far-distribution shifts involve more substantial differences, such as changes in satellite platforms, sensor types, or entirely different imagery domains, resulting in greater divergence from the training distribution. This variability necessitates tailored approaches for satellite-based machine learning models [26].
Despite the vital role of robust models in EO, few studies explore OOD detection in this domain [6, 3, 2, 32, 7]. These works often presume access to test-time distributions, adhere to closed-set assumptions, and/or degrade model performance on in-distribution tasks. This underscores the need for specialized OOD detection methods suited for large-scale EO deployment. A further challenge arises when applying geospatial models globally. Global geospatial models often suffer significant performance drops in low-data regions [1], such as sub-Saharan Africa [15]. These issues emphasize the need for global models equipped to detect OOD samples during inference.
@@
In this paper, we build on well-established finding that OOD samples trigger internal activation patterns that diverge from ID samples [30, 25]. Building on this insight and recognizing the challenges of deploying geospatial models at scale, we introduce the following key contributions:
  • We propose TARDIS (Test-time Addressing of Distribution Shifts at Scale), a lightweight post-hoc OOD detection method that preserves model performance and operates without access to labeled data from unknown test-time shifts.
  • We evaluate TARDIS on EuroSAT (patch classification) and xBD (semantic segmentation) under near-distribution settings, analyzing covariate (geographical, temporal, environmental) and semantic (withheld class) shifts across 17 experimental setups. We also compare its performance against existing post-hoc OOD detection methods.
  • We show that TARDIS scales effectively, providing actionable insights into the robustness and trustworthiness of geospatial models in real-world scenarios.
@@ 2 Related Works
Our method lies at the intersection of two notable categories of OOD detection methods: scoring functions and activation manipulation. Scoring functions assign a numerical score to each input, reflecting its alignment with ID samples based on the model’s output. Maximum Softmax Probability (MSP) [13] detects OOD samples by assessing the softmax confidence score, assuming that low confidence indicates OOD samples. The Out-of-DIstribution detector for Neural networks (ODIN) [20] enhances MSP by applying input perturbations and temperature scaling to improve ID-OOD separation. The Mahalanobis score [19] calculates the distance between input features and class means in feature space, flagging inputs far from these means as OOD. The energy score [22] evaluates the model’s energy function to assess the likelihood of an input belonging to the ID distribution. Another category of OOD detection methods manipulates the internal activations of a pre-trained model to improve detection performance. ReAct [30] identifies differences in activation patterns of the penultimate layer between ID and OOD samples, enhancing separation by clipping activations at an upper limit. DICE [29] applies weight sparsification on a specific layer to further distinguish ID from OOD data and, when combined with ReAct, can enhance detection performance. Activation Shaping [4] prunes a portion of an input sample’s activations and slightly adjusts the remaining activations; when paired with the energy score, this approach has been shown to outperform contemporary OOD methods. Neuron Activation Patterns (NAP) [25] extracts, downsamples, and binarizes activation patterns from convolutional layers, computing the smallest Hamming distance between binarized test and training patterns. This distance provides a measure of model uncertainty, aiding in OOD detection. Both scoring functions and activation-based strategies face notable challenges in dynamic, real-world environments. Scoring functions, while maintaining ID accuracy, struggle with OOD detection due to their reliance on static data distributions, an assumption rarely valid in practice. Activation-based methods, though effective, often compromise ID task performance and rely on hyperparameters tuned with hypothetical or substitute OOD samples. This reliance limits their effectiveness under evolving test-time conditions and shifting distributions.
Another line of OOD detection research uses clustering in feature space to separate ID and OOD samples. These methods exploit structural patterns in learned representations rather than direct scoring or activation manipulation. Gulati et al. [8] improve over $k$-means by using non-negative kernel regression for soft clustering. Sinhamahapatra et al. [27] explore clustering in self-supervised latent spaces, showing its potential in low-label settings. Closely related in spirit, NG-Mix [5] proposes pseudo-anomaly generation for semi-supervised anomaly detection.
@@
The task of OOD detection is inherently challenging, and the EO domain is no exception due to the diverse and heterogeneous nature of satellite imagery. Despite the critical importance of OOD detection in EO, this area remains relatively underexplored, with only a limited number of studies addressing the issue. One such study develops a Dirichlet Prior Network to quantify distributional uncertainty in deep learning models for satellite image analysis [6]. This method assumes Dirichlet distributions for ID samples and employs various setups where classes, color channels, and environmental features are alternately treated as ID or OOD. Another study frames anomaly detection as a cumulative open-set detection and location separation task [3]. This approach also uses a Dirichlet prior, like the first study, to expand the representation space between ID (normal) and OOD (anomalous) samples by predicting the anticipated categorical distribution. The method suits scenarios where pre-event images are unavailable, subject to radiation differences, or not recent enough to aid detection.
In this paper, we do not aim to optimize OOD detection performance but propose a method suited for real-world applications where the distribution is unknown, maintaining ID task performance is critical, and expensive methods are impractical. Our method detects shifts within the same satellite sources. This type of near-distribution shift presents a greater challenge than far-distribution shifts, such as those between natural and satellite imagery, as the differences are more subtle and harder to disentangle. By tackling these nuanced shifts, our method makes OOD detection more practical for real-world geospatial applications.
@@ 3 Problem Formulation
4 Method
Training a classifier to distinguish in-distribution (ID) from out-of-distribution (OOD) samples typically requires labeled ID and OOD data, which is impractical in real-world settings. This motivates generating surrogate labels for samples with unknown distributions, referred to as WILD samples. Our method, illustrated in Figure 1, operates with a pre-trained model, known ID samples, and WILD samples. Using known ID samples, it assigns surrogate ID and OOD labels to WILD samples, which are then used to train a binary distribution classifier. TARDIS involves training a binary classifier to discriminate between ID and OOD samples based on internal activations, as shown in Figure 2. Given $M$ ID and $N$ WILD samples, we pass both through the frozen model and extract activations from a specified layer. After pooling for spatial downsampling, we obtain one-dimensional feature vectors, combine them, and assign corresponding ID or WILD labels. We shuffle this set and reserve 30% for validation. On the remaining 70%, we run $k$-means clustering to assign surrogate ID/OOD labels to WILD samples based on distance to ID samples. Using this labeled set, we train a binary classifier $g(\mathbf{z};\theta)$ to map activations to distribution labels.
At deployment, we extract activations from a sample of unknown distribution and predict its label as $\hat{y}=g\left(\text{Downsample}\left(f(\mathbf{x})\right);\theta\right)$. The output can be interpreted as a probability of domain shift or thresholded for binary classification (00: ID, $1$: OOD).
Surrogate Label Assignment. The core novelty of TARDIS lies in generating surrogate labels for WILD samples based on their proximity to known ID samples in feature space. We assume that ID and OOD features are separable enough for clustering to be effective. Specifically, we cluster the combined ID+WILD feature space into $k$ groups (Step 3 in Figure 2). For each cluster, if the proportion of ID samples is greater than or equal to the threshold $T$, all samples in that cluster receive label 0 (ID); otherwise, label 1 (OOD). Distance to ID samples thus serves as a proxy for assigning surrogate labels.
Given $k$, $T$, and the feature vectors, we perform clustering and assign surrogate labels accordingly. To select the optimal parameters, we minimize the composite objective:
$(k^{*},T^{*})=\arg\min_{k,T}\quad H(S)+P_{\text{mis-ID}}-P_{\text{corr-ID}}$
where $H(S)$ is the average cluster entropy (encouraging homogeneity), $P_{\text{mis-ID}}$ is the proportion of ID samples misclassified as OOD, and $P_{\text{corr-ID}}$ is the proportion of correctly classified ID samples. This balances three goals: coherent clusters, low ID misclassification, and high correct ID classification, promoting better surrogate label quality.
@@ 5 Experimental Setup
In this section, we describe experimental setups to evaluate TARDIS under controlled distribution shifts, mimicking real-world geospatial deployment challenges. We use the EuroSAT [12] dataset for patch-level classification and the xBD [10] dataset for semantic segmentation. EuroSAT contains 27,000 labeled images across ten land use/land cover classes. xBD includes pre/post-disaster imagery labeled for building damage assessment, covering hurricanes, floods, wildfires, and more. For our purposes, we formulate xBD as a building footprint segmentation task with two classes: building footprint and background. These datasets offer diverse evaluation conditions in terms of volume, task, sensor, spatial/temporal coverage, and geolocation, making them suitable test-beds for our OOD method.
@@
To simulate controlled distribution shifts, we adjust the train, validation, and test splits in both datasets so that the model, trained on the modified training set, encounters realistic shifts during testing. We evaluate two shift types: covariate shift, where input distribution changes between train and test, and semantic shift, where unseen classes appear at test time. Covariate shifts include spatial (different locations), temporal (different time periods), and environmental (different landcover/disaster types) variance. For semantic shift, we train on nine of ten classes and test on the held-out class, rotating through all classes. Table 1 summarizes these setups.
These setups assume access to clear ID (train) and OOD (test) labels—an unrealistic assumption in real-world deployments. To address this, we evaluate our method in two ways. First, we use known ID and OOD labels to directly train and evaluate the binary classifier $g$, referred to as $g_{\text{oracle}}$. Second, we treat test-time OOD labels as unknown (i.e., WILD) and apply our surrogate label assignment to generate surrogate ID and OOD labels. This version, trained and evaluated using surrogate labels, is denoted as $g^{*}$. We consider $g_{\text{oracle}}$ as an upper bound on performance, since it uses ground-truth labels, while $g^{*}$ relies on inferred surrogates.